evaluate.simplify your answer write an exact answer using i as needed
i^32

Answer by Edwin McCravy(20065)


All you need to know is that 

i2 = -1
(-1)EVEN = 1
(-1)ODD = -1


i32 = i2*16 = (i2)16 = (-1)16 = 1

Another example with an even power of i:

i58 = i2*29 = (i2)29 = (-1)29 = -1

Another example with an odd power of i:

i75 = i74+1 = i74i1 = i2*37i = (i2)37i = (-1)37i = -1i = -i

Another example with an odd power of i:

i201 = i200+1 = i200i1 = i2*100i = (i2)100i = (-1)100i = 1i = i

Edwin